OpenStudy (dawnr):

V - domain in 1st quadrant limited with flat : x/2 + y/2 + z = 1

OpenStudy (dawnr):

I need to find V or volume

OpenStudy (dawnr):

which method do i use here?

OpenStudy (adrimit):

multi-integral

OpenStudy (adrimit):

in this case a triple integral

OpenStudy (adrimit):

what are the boundaries of your curve?

OpenStudy (dawnr):

i don't know how to find the boundaries

OpenStudy (adrimit):

ok in order to integrate something, you must first know its boundaries. do you know how to compute a single integral?

OpenStudy (dawnr):

yes..when i have it's boundaries

OpenStudy (adrimit):

find \(\int x\;dx\)

OpenStudy (dawnr):

here i tried using jacobian and changing x/2 , y/2 and z to u , v and w but i am not quite sure if i got the calculation right for the matrix

OpenStudy (adrimit):

you do not need a Jacobian here; only the boundaries.

OpenStudy (dawnr):

but how do i get them? what am i supposed to do with the given equation?

OpenStudy (dawnr):

integral of xdx is x^2 / 2

OpenStudy (adrimit):

recall that \(\frac{1}{2}x+\frac{1}{2}y+z=1\) is a plane

OpenStudy (adrimit):

when you integrate some curve of one independent variable, you add a set of rectangles

OpenStudy (dawnr):

so do i use the y = ax + b ?

OpenStudy (adrimit):

no, you must find where the plane \(\frac{1}{2}x+\frac{1}{2}y+z=1\) intersects the xy-plane, yz-plane and zx-plane.

OpenStudy (dawnr):

okay, how do i do that?

OpenStudy (adrimit):

ok, let us first observe the order of the differentials (dx dy dz)

OpenStudy (adrimit):

the limits of the first integral are 0 to the boundary of the given plane and the yz-plane

OpenStudy (adrimit):

i.e. where they intersect

OpenStudy (dawnr):

so where do they intersect? do i use any values for x and y or is there a way to calculate this?

OpenStudy (adrimit):

what is the x coordinate of any point on the yz-plane?

OpenStudy (michele_laino):

hint: the situation can be represented by this drawing: